# Pi in the Sky: Neutron Stars with Exceptionally Light QCD Axions

### v1

@@ 3.4 EOS of axion condensed matter
Figure 6 shows the RMFT EOS for charge neutral, $\beta$-equilibrated matter for several choices of $\varepsilon$ in the three scenarios discussed. We focus on the crust and outer regions of the core as those are the densities at which the effects of axion condensation are well understood. The solid red curve corresponds to the normal EOS without axions. For comparison, the black dotted curve shows the BPS+NV crust EOS BPS1971 ; NegeleVautherin1973 .
@@
We first focus on the left panel of Fig. 6. For $\varepsilon=10^{-6}$, the axion field itself contributes very little to the energy density and pressure at the densities pictured and Scenario A looks very similar to the case with no axions. At pressures of order $\varepsilon m_{\pi}^{2}f_{\pi}^{2}\simeq 10\times\varepsilon\,{\rm MeV/fm}^{3}$ the curve for Scenario A will diverge from the EOS without axions and go off the bottom of the plot as in the center panel. This occurs because zero pressure in the axion condensed phase occurs when the pressure of nuclear matter exactly cancels the negative axion pressure, corresponding to a much larger baryon density than normally found in nuclear matter at zero pressure. The point where the curves for Scenario B and C in the left panel diverge from Scenario A and the EOS without axions corresponds to the neutron drip phase transition for Scenario A and the EOS without axions. When $K_{\sigma}$ is large enough that the inner crust is absent, the pressure is larger (the EOS is “stiffer”) at energy densities where ordinarily dripped neutrons would be present because there are fewer degrees of freedom in the system, resulting in the lepton density being larger. At the crust-core transition, however, there is a large jump in energy density from the mixed phase to bulk nuclear matter, resulting in a lower pressure (the EOS is “softer”) in the outermost part of the core.
@@
The center and right panels of Fig. 6 can be understood in terms of the basic trends in the left panel. For $\varepsilon=10^{-3}$, the outermost regions of the star are still axion condensed and the only difference between the center and left panels is the much larger energy density and negative pressure from the axion field itself. This results in the equation of state going to zero pressure at large energy density because the nuclear matter at zero net pressure must contribute enough pressure to cancel the much larger axion pressure. In the right panel, Scenario B and C remain axion condensed down to zero pressure but the much larger value of $\varepsilon$ means that this occurs before the crust-core phase transition and the jump in energy density seen at lower $\varepsilon$ is not present. In Scenario A in the right panel, axion condensation is not favored at the lowest densities and there is a phase transition from axion condensed matter to normal matter, producing a discontinuity in the energy density due to the much lower baryon density in the normal phase.
@@
Figure 6: RMFT EoS for various $\varepsilon$ and $K_{\sigma}$ using IUFSU parameter set. BPS+NV EoS for the crust is shown for comparison BPS1971 ; NegeleVautherin1973 .
4 Neutron stars with $\theta=\pi$ matter
@@
In this section, we describe the modification to the neutron star EOS taking into account the effects we discussed in Secs. 2 and 3, including the negative pressure at $\theta=\pi$ and modifications to the nucleon masses and interactions, as well as the axion field profile and domain wall pressure. Neutron star structure can be derived by combining gravity and these new effects that depend on the axion coupling and masses, parameterized by $\varepsilon$ and $m_{a}$. Qualitatively, for larger $m_{a}$ ($m_{a}\gg 1/R_{\rm NS}$), the axion field profile can be treated as a thin wall over which the gravitational potential does not change, and the gravity and axion effects can be accounted for separately. We will first discuss neutron star structure in the simpler case where this approximation is valid, before commenting on the criteria for the validity of the approximation and working out the more general cases.
4.1 Neutron star structure in axion condensed stars
@@
An ordinary neutron star has an outer crust composed of nuclei in a background of free electrons and an inner crust composed of nuclei in a background of free electrons and dripped neutrons, together having a thickness of around a kilometer (see the left panel of Fig. 7 for a sketch). The possibilities for the outer structure of a neutron star become more varied in the presence of axion condensation, depending on the value of $\varepsilon$ and any modifications to the nuclear force. In the simplest case, for $\varepsilon\gtrsim 0.1$, the center of the star is axion condensed with a transition to normal matter at some intermediate radius. That radius may be large enough that portions of the normal crust are missing with, e.g., a transition directly from bulk matter at $\theta=\pi$ to an outer crust at $\theta=0$ with possibly a thin inner crust in the domain wall region. For larger $\varepsilon$, the entire normal crust is present and only a portion of the bulk matter in the center of the star is axion condensed, with a domain wall region of thickness $\Delta R\sim 1/m_{a,\rm eff}$ in which the energy density changes rapidly but the pressure is nearly constant. This case is sketched in the right panel of Fig. 7.
@@
Figure 7: Layers of a normal neutron star (left) and a neutron star with an axion condensed inner core (right). The thickness of layers is not to scale.
@@
Figure 8: Layers of a neutron star with a crust only in the domain wall region (left) and one with an axion condensed crust (right) in Scenario A. Scenario B and C are the same, but with no dripped neutron phase. The thickness of layers is not to scale.
@@
Although our treatment of axion condensed matter does not apply at arbitrarily high density, we can still speculate about the effects of axion condensation on neutron star masses and radii. As can be seen in Fig. 6, at low density, the EOS is typically softer for $\theta=\pi$ and $\varepsilon$ not too small because of the contribution of the axion field. A naive extrapolation of nuclear matter to high density with only the axion-nucleon interaction we have considered thus far will eventually yield a stiffer EOS because of the lighter nucleon mass. This results in heavy neutron stars having larger radii and larger maximum mass. Figure 9 shows the relationship between mass and radius for neutron stars with various $\varepsilon$ within this simplistic picture of the high density EOS and treating the domain wall as thin ($1/m_{a}\ll r_{\rm NS})$.
@@
Counter-intuitively, larger $K_{\sigma}$ yields a larger NS radii for very small $\varepsilon$, pointing to a stiffer EOS. At high density, larger $K_{\sigma}$ results in a stiffer EOS because of greater reduction to the effective nucleon mass. At low density, larger $K_{\sigma}$ corresponds to nuclei with larger isospin asymmetry increasing the electron pressure once the neutron dripped phase is missing, again giving a stiffer EOS. Only in a range of energy densities just above the crust-core phase transition do Scenario B and C give a softer EOS than Scenario A with Scenario C having such a small region of softening that it has minimal effect on the mass-radius curve and radii are larger for all masses.
@@
Figure 9: Mass-radius relationship for neutron stars with various $\varepsilon$ and $\theta$ in RMFT with only the leading interaction between nucleons and axions (introducing unknown uncertainties for $M\gtrsim 1.4\,M_{\odot}$) and neglecting the thickness of the domain wall.
@@
For $\varepsilon$ where the lowest density parts of the crust are missing, low mass stars may have very small radii, with a mass-radius relation akin to strange quark stars. Unlike a strange quark star, however, neutron stars with a condensed axion field lack a confining potential at their surface, and as a result have a thin crust of matter at intermediate $\theta$ in the domain wall region. Since the density and chemical potential smoothly go to zero on a length scale determined by the axion field profile, unusual surface properties of strange quark stars (see, e.g., Ref. StrangeStars for a detailed discussion) are not present in axion condensed stars.
@@
Since our treatment of the interaction between axions and nucleons is only strictly accurate at densities around or below nuclear density, predictions from our calculation of the mass-radius relationship should be considered speculative for $M\gtrsim 1.4\,M_{\odot}$. As such, the constraints we draw will all be based on the structure of the crust in axion condensed stars. We will use masses and radii of $1.4\,M_{\odot}$ stars calculated with the uncertain high density equation of state as a starting point to calculate the crust and domain wall, but these will only have an effect by altering the gravitational potential and explicit dependence of gradients on the radius in spherical coordinates. Since the inclusion of axion condensation changes the radius of a $1.4\,M_{\odot}$ star by up to approximately a kilometer, and the uncertainties from nuclear physics are ±$\pm 1-3\,{\rm km}$ (cf. Fig. 3 in Ref. PRC_limitingMR ), this is a sub-leading source of uncertainty.
@@
Figure 10: Crust mass and crust thickness of a $1.4\,M_{\odot}$ neutron star as a function of $\varepsilon$ for RMFT EOS, normalized to 1 for star without axions.
@@
Focusing on the crust, Fig. 10 shows the total thickness and mass of the crust and thickness and mass of the outer crust alone of a $1.4\,M_{\odot}$ neutron star as a function of $\varepsilon$, normalized to the value for a star without axions. In Scenario A, the stars at very small $\varepsilon$ have nearly the entire crust and look very similar to a star without axions, though the envelope (the region of the crust with $\rho<10^{10}\,{\rm g/cm}^{3}$) is still diminished, providing a constraint from the cooling of isolated neutron stars, as will be discussed in Sec. 6.2. In Scenarios B and C, there is less overall mass in the crust because of the lack of dripped neutrons. This will reduce the moment of inertia in the crust and can be constrained with pulsar glitches, as will be discussed in Sec. 6.3. In all three scenarios, when $10^{-4}\lesssim\varepsilon\lesssim 0.1$ the overall thickness of the crust is diminished. Observations of the thermal relaxation timescale of neutron stars in x-ray binaries puts a lower limit on the necessary thickness of the crust and will be discussed in Sec. 6.1.
4.2 Axion domain wall
@@
An axion domain wall separates regions with densities that are larger and smaller than $n_{B}\sim\varepsilon m_{\pi}^{2}f_{\pi}^{2}/\sigma_{N}$. This profile, in general, should be solved together with the density and pressure profile of the neutron star. In the limit where $m_{a}^{\rm QCD}>m_{a}>1/r_{\rm NS}$, the axion profile is a thin domain wall, located at a radius $r_{a}$ from the center of the neutron star. The tension $\sigma_{a}$ of this field profile wrapping the nuclear matter inside the neutron star exerts a pressure of $\sigma_{a}/2r_{a}$, which shall be considered as the minimal pressure of the matter inside the domain wall.
@@
We will consider two limits to solve the axion field profile, extending the formalism described in  2.3 to include nuclear interactions. Firstly, we will consider an axion profile in a smooth background distribution of nuclear matter that is changing slowly over length scales that are much longer than $1/m_{a}$. In this case, we can solve the axion field profile in PNM as
@@
$\begin{split}\nabla^{2}a&=\left(m_{a}^{2}-\frac{\sigma_{N}(1+\delta^{\rm(loc)})n_{n}(r)}{f_{a}^{2}}\right)f_{a}\,\sin[a/f_{a}]\\ &\simeq\frac{\sigma_{N}(1+\delta^{\rm(loc)})}{f_{a}}\left.\frac{\mathrm{d}n_{n}(r)}{\mathrm{d}r}\right|_{r=r_{a}}(r-r_{a})\,\sin[a/f_{a}].\end{split}$ (23)
@@
where $\delta^{\rm(loc)}\propto K_{\sigma}n_{B}/\sigma_{N}m_{\sigma}^{2}$ is an order one quantity that encodes the $m_{\sigma}$ dependence of the effective axion mass. Details can be found in App. B.
@@
For axion profiles with thickness much smaller than $r_{a}$, at linear order in $a/f_{a}$, the solution of this equation is an Airy function, where the field profile changes over a length scale of $(m_{a}^{2}/r_{a})^{-1/3}$. The domain wall tension of this axion profile is $\sigma_{a}\leq m_{a}f_{a}^{2}$ and the resulting pressure difference due to this domain wall profile between inside and outside is smaller than the pressure difference due to the vacuum energy difference of $m_{a}^{2}f_{a}^{2}$.
@@
The smooth density profile would be a good approximation only for the largest $\varepsilon$ where axion domain wall is situated at close to nuclear densities. As we show in Sec. 3, for most of the $\varepsilon$, the domain wall is between regions with drastically different densities, and the density (and pressure) of the matter outside the domain wall is effectively zero. In this second case, inside the domain wall, the density $n_{\rm in}$ is such that the degeneracy pressure is
@@
This fails in the case where the domain wall is very near the center of the star and $M_{a}$ is small. In practice, this is a challenging constraint to implement since $m_{a,\rm eff}$ changes sign in the vicinity of the domain wall. In the core where the baryon density changes more slowly, this leads to a very extended domain wall even for small axion masses, with the extent of the domain wall becoming comparable to the gravitational scale for $m_{a}\lesssim 10^{-9}\,\mbox{eV}$. For the sake of calculating the contribution of the domain wall to the crust, we limit ourselves to axion masses $m_{a}\gtrsim 10^{-9}\,\mbox{eV}$ when the domain wall is inside the star. This does not affect the final constraint we determine.
@@
where $n^{(s)}_{i}=\partial E/\partial m_{i}$ is the scalar density of baryon species $i$ which at leading order is given by $n^{(s)}_{i}=n_{i}-k_{Fi}^{5}/[10\pi^{2}m_{i}^{2}(\theta)]$, the spatial gradients include corrections from gravity, and the scalar densities are found at fixed baryon chemical potential. The quantity $\delta^{\rm(loc)}$ comes from the explicit $m_{\sigma}$-dependence of the RMFT Lagrangian and is approximately given by
$\delta^{\rm(loc)}\approx\frac{K_{\sigma}}{\chi\sigma_{N}}\left(\frac{g_{\sigma}}{m_{\sigma}}\right)^{2}(n_{p}^{(s)}+n_{n}^{(s)})$ (30)
where $g_{\sigma}$ is the coupling of the $\sigma$ meson to nucleons in the RMFT model and $\chi$ is the volume fraction of nuclei in the outer crust and one in the core. Exact expressions including higher order RMFT corrections and the more complicated expression for the inner crust can be found in App. B. This correction is order one for both Scenarios B and C in the outer crust and core and should not be neglected, but is suppressed in the inner crust. Note that this correction is due only to explicit $m_{\sigma}$ dependence of the RMFT Lagrangian; implicit dependence via the effect of $m_{\sigma}$ on the Fermi momenta of the nucleons and the mean field value of the $\sigma$ field vanishes since the mean field approximation will always give Fermi momenta and a mean field value for the $\sigma$ field that minimize the free energy.
@@
The differential equation given by Eq. (29) can be solved by a shooting procedure, varying the initial conditions at the inner edge of the domain wall until a solution is found that smoothly goes from $\theta=\pi$ to zero. In practice, solving this equation is greatly simplified by making a change of variables to confine $\theta$ to be between zero and $\pi$, $\theta(y)\equiv\pi/(1+e^{y})$ and solving the resulting equation for $y$. Numerical results from this calculation will be presented in Sec. 6.
Since we perform calculations in the domain wall region only in RMFT, the entire $\theta$-dependence in the domain wall is contained in the nucleon and meson masses and the energy of the axion field itself. There are additional terms in the nuclear Lagrangian that contribute only when $\theta$ is at some intermediate value that violate parity and a detailed calculation of the domain wall from first principles should include these terms that we neglect. For further discussion, see App. A.
5 Finite size objects at zero pressure
An object comprised of nuclear matter at $\theta=\pi$ can be stabilized by the negative pressure $m_{a}^{2}f_{a}^{2}$, rather than gravity. These objects, which we call $\pi$-balls, are held together by the attractive axion potential as long as the baryon chemical potential is too small to allow nucleons to escape (see also  Witten:1984rs ; Zhitnitsky:2002qa ). Given the constraint from white dwarf stability of $\varepsilon>2\times 10^{-7}$, a zero-pressure object balanced by electron degeneracy pressure cannot exist. If a neutron star has a central density large enough to be unstable against axion condensation, it is possible that when such a neutron star is disrupted in a binary neutron star merger, $\pi$-balls may be part of the ejecta. Having an axion field surrounding them, these objects may produce a unique electromagnetic signature. Such questions of production and observational signatures of $\pi$-balls will be explored in the future; in this work, we merely comment on their basic properties.
Figure 11: Minimum stable radius of a $\pi$-ball made of SNM and PNM. Thin lines: $m_{a}=10^{-9}\,\mbox{eV}$, thick lines: $m_{a}=10^{-6}\,\mbox{eV}$. The dotted black lines show lines of $R=\pi/m_{a}$ for the two mass cases.

### v2

@@ III.4 EOS of axion condensed matter
Figure 6 shows the RMFT EOS for charge neutral, $\beta$-equilibrated matter for several choices of $\varepsilon$ in the three scenarios discussed. We focus on the crust and outer regions of the core as those are the densities at which the effects of axion condensation are well understood. The solid red curve corresponds to the normal EOS without axions. For comparison, the black dotted curve shows the BPS+NV crust EOS [77, 78].
@@
We first focus on the left panel of Fig. 6. For $\varepsilon=10^{-6}$, the axion field itself contributes very little to the energy density and pressure at the densities pictured, and Scenario A looks very similar to the case with no axions. At pressures of order $\varepsilon m_{\pi}^{2}f_{\pi}^{2}\simeq 10\times\varepsilon\,{\rm MeV/fm}^{3}$ the curve for Scenario A will diverge from the EOS without axions and go off the bottom of the plot as in the center panel. This occurs because zero pressure in the axion condensed phase occurs when the pressure of nuclear matter exactly cancels the negative axion pressure, corresponding to a much larger baryon density than normally found in nuclear matter at zero pressure. The point where the curves for Scenario B and C in the left panel diverge from Scenario A and the EOS without axions corresponds to the neutron drip phase transition for Scenario A and the EOS without axions. When $K_{\sigma}$ is large enough that the inner crust is absent, the pressure is larger (the EOS is “stiffer”) at energy densities where ordinarily dripped neutrons would be present because there are fewer degrees of freedom in the system, resulting in the lepton density being larger. At the crust-core transition, however, there is a large jump in energy density from the mixed phase to bulk nuclear matter, resulting in a lower pressure (the EOS is “softer”) in the outermost part of the core.
@@
The center and right panels of Fig. 6 can be understood in terms of the basic trends in the left panel. For $\varepsilon=10^{-3}$, the outermost regions of the star are still axion condensed, and the only difference between the center and left panels is the much larger energy density and negative pressure from the axion field itself. This results in the equation of state going to zero pressure at large energy density because the nuclear matter at zero net pressure must contribute enough pressure to cancel the much larger axion pressure. In the right panel, Scenario B and C remain axion condensed down to zero pressure but the much larger value of $\varepsilon$ means that this occurs before the crust-core phase transition and the jump in energy density seen at lower $\varepsilon$ is not present. In Scenario A in the right panel, axion condensation is not favored at the lowest densities and there is a phase transition from axion condensed matter to normal matter, producing a discontinuity in the energy density due to the much lower baryon density in the normal phase.
@@
Figure 6: The RMFT EoS for three example values of $\varepsilon$ and the three $K_{\sigma}$ scenarios using the IUFSU parameter set. The standard no axion RMFT EoS and the BPS+NV EoS for the crust are shown for comparison [77, 78]. See section III.4 for details.
IV Neutron stars with $\theta=\pi$ matter
@@
In this section, we describe the modification to the neutron star EOS taking into account the effects we discussed in Secs. II and III, including the negative pressure at $\theta=\pi$ and modifications to the nucleon masses and interactions, as well as the axion field profile and domain wall pressure. Neutron star structure can be derived by combining gravity and these new effects that depend on the axion coupling and masses, parameterized by $\varepsilon$ and $m_{a}$. Qualitatively, for larger $m_{a}$ ($m_{a}\gg 1/R_{\rm NS}$), the axion field profile can be treated as a thin wall over which the gravitational potential does not change, and the gravity and axion effects can be accounted for separately. We will first discuss neutron star structure in the simpler case where this approximation is valid, before commenting on the criteria for the validity of the approximation and working out the more general cases.
IV.1 Neutron star structure in axion condensed stars
@@
An ordinary neutron star has an outer crust composed of nuclei in a background of free electrons and an inner crust composed of nuclei in a background of free electrons and dripped neutrons, together having a thickness of around a kilometer (see the left panel of Fig. 7 for a sketch). The possibilities for the outer structure of a neutron star become more varied in the presence of axion condensation, depending on the value of $\varepsilon$ and any modifications to the nuclear force. In the simplest case, for $\varepsilon\gtrsim 0.1$, the center of the star is axion condensed with a transition to normal matter at some intermediate radius. That radius may be large enough that portions of the normal crust are missing with, e.g., a transition directly from bulk matter at $\theta=\pi$ to an outer crust at $\theta=0$ with possibly a thin inner crust in the domain wall region. For larger $\varepsilon$, the entire normal crust is present and only a portion of the bulk matter in the center of the star is axion condensed, with a domain wall region of thickness $\Delta r_{\rm DW}\sim 1/m_{a,\rm eff}$ in which the energy density changes rapidly but the pressure is nearly constant. This case is sketched in the right panel of Fig. 7.
@@
Figure 7: Layers of a normal neutron star (left) and a neutron star with an axion condensed inner core (right). The thickness of layers is not to scale. See section IV.1 for discussion.
@@
Figure 8: Layers of a neutron star with a crust only in the domain wall region (left) and one with an axion condensed crust (right) in Scenario A. Scenario B and C are the same, but with no dripped neutron phase. The thickness of layers is not to scale. See section IV.1 for discussion.
@@
Although our treatment of axion condensed matter does not apply at arbitrarily high density, we can still speculate about the effects of axion condensation on neutron star masses and radii. A similar calculation was performed in Ref. [45] for a non-interacting gas of neutrons. As can be seen in Fig. 6, at low density, the EOS is typically softer for $\theta=\pi$ and $\varepsilon$ not too small because of the contribution of the axion field. A naive extrapolation of nuclear matter to high density with only the axion-nucleon interaction we have considered thus far will eventually yield a stiffer EOS because of the lighter nucleon mass. This results in heavy neutron stars having larger radii and larger maximum mass. Figure 9 shows the relationship between mass and radius for neutron stars with various $\varepsilon$ within this simplistic picture of the high density EOS and treating the domain wall as thin ($1/m_{a}\ll r_{\rm NS})$.
@@
Counter-intuitively, larger $K_{\sigma}$ yields larger NS radii for very small $\varepsilon$, pointing to a stiffer EOS. At high density, larger $K_{\sigma}$ results in a stiffer EOS because of greater reduction to the effective nucleon mass. At low density, larger $K_{\sigma}$ corresponds to nuclei with larger isospin asymmetry increasing the electron pressure once the neutron dripped phase is missing, again giving a stiffer EOS. Only in a range of energy densities just above the crust-core phase transition do Scenario B and C give a softer EOS than Scenario A with Scenario C having such a small region of softening that it has a minimal effect on the mass-radius curve and radii are larger for all masses.
@@
Figure 9: Mass-radius relationship for neutron stars with three fiducial $\varepsilon$ values and $K_{\sigma}$ in RMFT. We consider only the leading interaction between nucleons and axions (which leads to unknown uncertainties for $M\gtrsim 1.4\,M_{\odot}$) and neglect the thickness of the domain wall.
@@
For $\varepsilon$ where the lowest density parts of the crust are missing, low mass stars may have very small radii, with a mass-radius relation akin to strange quark stars. Unlike a strange quark star, however, neutron stars with a condensed axion field lack a confining potential at their surface, and as a result have a thin crust of matter at intermediate $\theta$ in the domain wall region. Since the density and chemical potential smoothly go to zero on a length scale determined by the axion field profile, unusual surface properties of strange quark stars (see, e.g., Ref. [79] for a detailed discussion) are not present in axion condensed stars.
@@
Since our treatment of the interaction between axions and nucleons is only strictly accurate at densities around or below nuclear density, predictions from our calculation of the mass-radius relationship should be considered speculative for $M\gtrsim 1.4\,M_{\odot}$. As such, the constraints we draw will all be based on the structure of the crust in axion condensed stars. We will use masses and radii of $1.4\,M_{\odot}$ stars calculated with the uncertain high density equation of state as a starting point to calculate the crust and domain wall, but these will only have an effect by altering the gravitational potential and explicit dependence of gradients on the radius in spherical coordinates. Since the inclusion of axion condensation changes the radius of a $1.4\,M_{\odot}$ star by up to approximately a kilometer, and the uncertainties from nuclear physics are ±$\pm 1$-$3\,{\rm km}$ (cf. Fig. 3 in Ref. [80]), this is a sub-leading source of uncertainty.
@@
Figure 10: Crust and outer cruss mass and thickness of a $1.4\,M_{\odot}$ neutron star as a function of $\varepsilon$ for RMFT EOS. The three panels correspond to the three interaction scenarios (Table 1). The curves are normalized to $1$ for a star without axions.
@@
Focusing on the crust, Fig. 10 shows the total thickness and mass of the crust and thickness and mass of the outer crust alone of a $1.4\,M_{\odot}$ neutron star as a function of $\varepsilon$, normalized to the value for a star without axions. In Scenario A, the stars at very small $\varepsilon$ have nearly the entire crust and look very similar to a star without axions, though the envelope (the region of the crust with $\rho<10^{10}\,{\rm g/cm}^{3}$) is still diminished, providing a constraint from the cooling of isolated neutron stars, as will be discussed in Sec. VI.2. In Scenarios B and C, there is less overall mass in the crust because of the lack of dripped neutrons. This will reduce the moment of inertia in the crust and can be constrained with pulsar glitches, as will be discussed in Sec. VI.3. In all three scenarios, when $10^{-4}\lesssim\varepsilon\lesssim 0.1$ the overall thickness of the crust is diminished. Observations of the thermal relaxation timescale of neutron stars in x-ray binaries put a lower limit on the necessary thickness of the crust and will be discussed in Sec. VI.1.
IV.2 Axion domain wall
@@
An axion domain wall separates regions with densities that are larger and smaller than $n_{B}\sim\varepsilon m_{\pi}^{2}f_{\pi}^{2}/\sigma_{N}$. This profile, in general, should be solved together with the density and pressure profile of the neutron star. In the limit where $m_{a}^{\rm QCD}>m_{a}>1/r_{\rm NS}$, the axion profile is a thin domain wall, located at a radius $r_{a}$ from the center of the neutron star. The tension $\sigma_{a}$ of this field profile wrapping the nuclear matter inside the neutron star exerts a pressure of $\sigma_{a}/2r_{a}$, the minimal pressure of the matter inside the domain wall.
@@
We solve the axion field profile in two regimes, extending the formalism described in section II.3 to include nuclear interactions. First, consider the limit where an axion domain wall is in a smooth background of nuclear matter which is changing slowly over length scales much longer than $1/m_{a}$. In this case, we can solve the axion field profile in PNM as
@@
$\begin{split}\nabla^{2}a&=\left(m_{a}^{2}-\frac{\sigma_{N}(1+\delta^{\rm(loc)})n_{n}(r)}{f_{a}^{2}}\right)f_{a}\,\sin[a/f_{a}]\\ &\simeq\frac{\sigma_{N}(1+\delta^{\rm(loc)})}{f_{a}}\left.\frac{\mathrm{d}n_{n}(r)}{\mathrm{d}r}\right|_{r=r_{a}}(r-r_{a})\,\sin[a/f_{a}],\end{split}$ (23)
@@
where $\delta^{\rm(loc)}\propto K_{\sigma}n_{B}/\sigma_{N}m_{\sigma}^{2}$ is an order-one quantity that encodes the $m_{\sigma}$ dependence of the effective axion mass. Details can be found in App. B.
@@
For axion profiles with thickness much smaller than $r_{a}$, at linear order in $a/f_{a}$, the solution of this equation is an Airy function, where the field profile changes over a length scale of $(m_{a}^{2}/r_{a})^{-1/3}$. The domain wall tension of this axion profile is $\sigma_{a}\leq m_{a}f_{a}^{2}$ and the resulting pressure difference due to this domain wall profile between inside and outside is smaller than that due to the vacuum energy difference of $m_{a}^{2}f_{a}^{2}$.
@@
The smooth density profile would be a good approximation only for the largest $\varepsilon$, where the axion domain wall is situated close to nuclear density. As we showed in Sec. III, for most of the $\varepsilon$ we consider, the domain wall is between regions with drastically different densities, and the density (and pressure) of the matter outside the domain wall is effectively zero. In this second case, inside the domain wall, the density $n_{\rm in}$ is such that the degeneracy pressure is
@@
This approximation fails in the case where the domain wall is very near the center of the star and $M_{a}$ is small. In practice, this is a challenging constraint to implement since $m_{a,\rm eff}$ changes sign in the vicinity of the domain wall. In the core where the baryon density changes more slowly, this leads to a very extended domain wall even for axion masses that are larger than $1/R_{\rm NS}$, with the extent of the domain wall becoming comparable to the gravitational scale for $m_{a}\lesssim 10^{-9}\,\mbox{eV}$. For the sake of calculating the contribution of the domain wall to the crust, we limit ourselves to axion masses $m_{a}\gtrsim 10^{-9}\,\mbox{eV}$ when the domain wall is inside the star. This does not affect the final constraint we determine.
@@
where $n^{(s)}_{i}=\partial E/\partial m_{i}$ is the scalar density of baryon species $i$ which at leading order in bulk matter is given by $n^{(s)}_{i}=n_{i}-k_{Fi}^{5}/[10\pi^{2}m_{i}^{2}(\theta)]$, the spatial gradients include corrections from gravity, and the scalar densities are found at fixed baryon chemical potential. The quantity $\delta^{\rm(loc)}$ comes from the explicit $m_{\sigma}$-dependence of the RMFT Lagrangian and is approximately given by
$\delta^{\rm(loc)}\approx\frac{K_{\sigma}}{\chi\sigma_{N}}\left(\frac{g_{\sigma}}{m_{\sigma}}\right)^{2}\left(n_{p}^{(s)}+n_{n}^{(s)}\right),$ (30)
where $g_{\sigma}$ is the coupling of the $\sigma$ meson to nucleons in the RMFT model and $\chi$ is equal to the volume fraction of nuclei in the outer crust and unity in the core. Exact expressions including higher-order RMFT corrections and the more complicated expression for the inner crust can be found in App. B. This correction is order one for both Scenarios B and C in the outer crust and core and should not be neglected, but is suppressed in the inner crust. Note that this correction is due only to explicit $m_{\sigma}$ dependence of the RMFT Lagrangian; implicit dependence via the effect of $m_{\sigma}$ on the Fermi momenta of the nucleons and the mean-field value of the $\sigma$ field vanishes since the mean-field approximation will always give Fermi momenta and a mean-field value for the $\sigma$ field that minimize the free energy.
@@
The differential equation given by Eq. (29) can be solved by a shooting procedure, varying the initial conditions at the inner edge of the domain wall until a solution is found that smoothly goes from $\theta=\pi$ to zero. In practice, solving this equation is greatly simplified by making a change of variables to confine $\theta$ to be between zero and $\pi$, $\theta(y)\equiv\pi/(1+e^{y})$ and solving the resulting equation for $y$. Numerical results from this calculation will be presented in Sec. VI.
Since we perform calculations in the domain wall region only in RMFT, the entire $\theta$-dependence in the domain wall is contained in the nucleon and meson masses and the energy of the axion field itself. There are additional terms in the nuclear Lagrangian that contribute only when $\theta$ is at some intermediate value that violates parity and a detailed calculation of the domain wall from first principles should include these terms that we neglect. For further discussion, see App. A.
V Finite size objects at zero pressure
As was first noted in Refs. [37, 45], an object comprised of nuclear matter at $\theta=\pi$ can be stabilized by the negative pressure $m_{a}^{2}f_{a}^{2}$, rather than gravity. These objects, which we call $\pi$-balls, are held together by the attractive axion potential as long as the baryon chemical potential is too small to allow nucleons to escape (see also  [81, 82]). Given the constraint from white dwarf stability of $\varepsilon>2\times 10^{-7}$, a zero-pressure object balanced by electron degeneracy pressure cannot exist. If a neutron star has a central density large enough to be unstable against axion condensation, it is possible that when such a neutron star is disrupted in a binary neutron star merger, $\pi$-balls may be part of the ejecta. Having an axion field surrounding them, these objects may produce a unique electromagnetic signature. Such questions of production and observational signatures of $\pi$-balls will be explored in the future; in this work, we merely comment on their basic properties.
Figure 11: Minimum stable radius of a $\pi$-ball made of SNM and PNM as a function of $\varepsilon$. Thin lines: $m_{a}=10^{-9}\,\mbox{eV}$, thick lines: $m_{a}=10^{-6}\,\mbox{eV}$. The dotted black lines show lines of $R=\pi/m_{a}$ for the two axion masses.